# Decoding Gray Matter: large-scale analysis of brain cell morphometry to inform microstructural modeling of diffusion MR signals

### v2

@@ Abstract
Diffusion-weighted MRI (dMRI) is a formidable technique for non-invasively characterizing brain microstructure. Biophysical modelling is often necessary to gain specificity to cellular structure.
However, designing sensible biophysical models and appropriate dMRI acquisitions is challenging, especially for gray matter (GM), as little is known about typical values of relevant features of braincell morphology contributing to dMRI signal. This work addressees this unmet need: We analysed about 3,500 cells from mouse, rat, monkey and human brains to determine statistical distributions of a set of morphological features (structural, topological and shape descriptors) relevant to GM microstructure modelling.
Illustrative examples demonstrate how this study can inform biophysical modelling and important implications for biophysical modelling of GM are discussed, enabling the design of more sensible models and suitable data acquisitions.
@@ 1 Introduction
Grey matter (GM) is composed of a range of cells, mainly differentiated into neuronal and glial cells, featuring a plethora of morphological characteristics. Neurons are fundamental functional units of the nervous system, specialized in the transmission and integration of electrical and chemical signals within the brain. Supporting the neurons are the glial cells (e.g. astrocites, microglia and olignodendracytes) that are crucial in maintaining the health and functionality of neurons.
First studied in depth by Ramon y Cajal in the late 19th century [1], neuronal morphology offers insights into the complex structure and function of the brain. Neurons come in various shapes and sizes, tailored for their respective functions [2]. Neurons can be classified according to their morphological, molecular, and physiological characteristics [3]. Both neurons and glia have similar structures, both having a central cell body (or soma) and branching projections, and show remarkable diversity and variability depending on the brain region [4, 5], e.g., from the complex branching patterns of Purkinje cells in the cerebellum to the more simple granule cells of the cerebral cortex.
@@
The brain contains about 86 billion neural111neural should not be confused with neuronal. Neural means relating to the nervous system (hence referring to any cell type in the brain: from neurons to glia), while neuronal means relating to neurons. cells [6]. Cortical GM is roughly constituted of 10–40% cell bodies (soma) of neural cells; 40–75% neurites: neuronal dendrites, short‐range intra‐cortical axons, the stems of long‐range axons extending into the WM and glial cell projections which intermingle with each other to form a dense and complex network; 15–30% highly tortuous extra-cellular space (ECS); and 1–5% vasculature [7, 8, 9]. In adults, the glia to neuron ratio is 1.32/1.40 for males/females respectively. The proportion of glial cells (by cell count) was estimated to be 77% oligodendrocytes, 17% astrocytes and 6% microglia [10, 11].
The ECS occupies a volume fraction of  15-30% in normal adult brain tissue, with a typical value of  20%, that falls to  5% during global ischemia (the expected state during classical fixation) [12]. The ECS has an average tortuosity (defined as the ratio between the true diffusion coefficient and the effective diffusivity of small molecules such as inulin and sucrose) of  2-3 [12], due to its labyrinthine porous matrix, the presence of long-chain macromolecules, transient trapping in dead-space microdomains, and transient physical-chemical interaction with the cellular membranes. The width of the ECS of the brain cortex is  40-120 nm [13], although this is less established because chemicals used in the processing of the tissue for the display introduce shrinkage effects that can range from 1-65% [14]. The average neuron-microvessel distance in brain GM is  20 $\mu$m [15].
@@
The morphology of neurons and glia is revealed through staining and microscopy techniques such as confocal or electron microscopy [16, 17, 18], which can image cellular structures with very high resolution (down to a few nanometers) [19]. However, currently there are no means to directly observe the cellular microstructure in vivo and without invasively using imaging techniques, such as Magnetic Resonance Imaging (MRI), as the cellular scale (on micrometers) lies beyond the resolution of clinical MRI, generally on the mm scale [20]. With many neurological conditions, including aging [21], altering the brain structure on this cellular scale, there is a strong incentive to develop means of revealing the morphology of neural cells, in-vivo.
@@
Given its sensitivity to the micrometer length scale, diffusion-weighted MRI (dMRI) is a promising technique to address the resolution limit of MRI and characterize the brain structure in vivo at the cellular scale, that is, the microstructure. However, this sensitivity is indirect and biophysical modelling of the dMRI signal in the brain tissue is essential to quantify histologically meaningful features of the cellular structure and gain specificity to their changes. To this end, the microstructure imaging paradigm has been introduced over a decade ago [14]: the approach voxel-wise fits a biophysical model to the set of signals obtained from images acquired with different sensitivities, yielding model parameter maps proxy of the corresponding microstructural features.
@@
Successful examples of the microstructure imaging paradigm include Diffusion Tensor Imaging (DTI) [22], the Neurite Orientation Dispersion and Density Imaging (NODDI) [23] and the White Matter Tract Integrity (WMTI) [24] to characterise the diffusion of water within white matter (WM), revealing insight into the structure of axonal bundle tracks and other anatomical features, such as axon diameter [25, 26], opening new and exciting opportunities in the field of neuroscience.
@@
Following the success of dMRI in WM, there has been a significant increase in interest in its application in GM, with the hopes of characterizing the cellular morphology within the brain in vivo. So far, dMRI has been used to differentiate different areas of cortical grey matter and reveal laminar structure within grey matter [27]. Great effort has been made to develop also models to better describe the diffusion signal within GM and reveal anatomical information about the microstructure, such as the Soma And Neurite Density Imaging (SANDI) [28] to characterize soma and neurite density, the Neurite EXchange Imaging (NEXI) [29] and the Standard Mmodel with EXchange (SMEX) [30] to characterize the water permeative exchange between neurites and extracellular space, and combinations of these, such as SANDI with exchange (SANDIx) [30].
@@
While substantial effort has been made to design, validate, and translate to clinics biophysical models for WM microstructure imaging [14, 31, 32], the GM counterpart is lagging. This is due to the greater complexity of the tissue, which renders the design of biophysical models for GM microstructure imaging more challenging [33].
To design sensible biophysical models and appropriate dMRI acquisitions for GM, it is critical to identify which relevant microstructural features of the complex GM tissue can be measured, how they can be estimated accurately and how such estimates can be validated. It is also very helpful to identify the typical values of such relevant features. With cellular morphology varying by cell type as well as brain region, there is a necessity for a thorough analysis of characteristics based on these criteria. Whilst substantial effort has been put towards the fine characterisation of WM axonal morphology there is yet to be the same level of in-depth analysis dedicated to GM cellular morphology.
@@
Here we aim to correct this imbalance with a comprehensive analysis of GM cellular morphology, looking at structural and topological morphology, and shape descriptors of over 3,500 real three-dimensional reconstructions from mouse, rat, monkey, and human brain cortex.
@@
The paper is organized as following. We first provide quantitative information on the anatomy of brain GM tissue at the cellular scale, which is the primary target for microstructure imaging through dMRI. We characterize the morphology of neural cells using structural, topological and shape descriptors. We then review the range of dMRI measurements and biophysical models available to probe this anatomy and highilight limitations and caveats, ultimately providing guidelines on how to model dMRI signals in GM.
2 Methods
@@ 2.1 Microscopy Dataset
In order to measure characteristic features for specific cell types, the open access repository Neuromorpho.org [34] was used, for comprehensive range of cellular reconstructions available. We downloaded and analysed three dimensional reconstructions of 3,598 brain cells, from mouse, rat, monkey, and human, in the form of SWC files. The ”SWC” encodes for the last names of its initial designers Ed Stockley, Howard Wheal, and Robert Cannon and is an ASCII text-based file that describes three-dimensional neuronal or glial morphology. The SWC file defines a set of labelled nodes connected by edges characterizing the three-dimensional structure of each cell Fig.1.
The cells acquired were representative of eight cell types: microglia, astrocyte, pyramidal, granule, purkinje, glutamatergic, gabaergic, and basket cells, from mouse/rat (N=1,358), monkey (N=948), and human (N=1,292). Although the NeuroMorpho database contains over 270,000 cellular reconstructions, only 4,278 satisfied our inclusion criteria: healthy controls; having complete reconstruction of all the dendrites; containing full information about diameters and angles; and being three-dimensional reconstructions. Of these, 680 did not pass our visual inspection on the quality of the reconstruction and were excluded. Our quality assessment criteria include: consistent estimates of dendritic diameters (e.g., most of the 680 reconstructions had a fixed nominal diameter for all the branches instead of the real one); continuity of the cellular processes; minimal reconstruction artifacts (e.g., most of the 680 reconstructions had artifactual shrinkage in the Z direction).
To reveal the skeletal structure of the cells, dendritic spines (small protrusions situated on neuronal dendrites, providing synaptic connections) were not included in the cellular reconstructions analysed to avoid them to bias the statistics of the overall cellular morphology, e.g. drastically shortening the branch length and increasing the branch order.
The cellular reconstructions were analysed in matlab (Mathsworks) using custom scripts, exploiting functions from validated suites (TREES [35], Blender [36], ToolboxGraph [37]). All the codes and SWC files used in this work will be made publicly available on GitHub upon publication.
@@ 2.2 Structural descriptors
The structural analysis describes the constituent parts of the cell, providing information about the general properties of the cell, the soma and the cellular projections. These features provide vital information about the fundamental skeleton of the cell such as effective soma radius and the branch angle formed between two daughter branches.
For the structural analysis, features or descriptors, determined to be crucial to microstructure modelling based on current literature [38, 28, 39, 40, 41, 42, 43, 44, 45, 46, 47], were estimated from the real data. This set of features enables us to understand better how each basic constituent of the brain cell morphology can impact the diffusion of molecules within the intra-cellular space. For example, the size of the soma and the cellular projections can inform on the characteristic length scales of intra-cellular restrictions; cellular projections’ branching, tortuosity, undulation and calibre coefficient of variation can inform on time dependence, and the soma and projections surface-to-volume ratios can inform on exchange.
The features of the soma and cellular projections are very different, so our analysis reflects that by dividing the structural descriptors into appropriate categories. These categories are: soma (the characteristics defining the cell body), projections (the set of characteristics defining the cellular projections’ structure as interconnected branches) and general (describing the general cellular characteristics).
Soma
  • $R_{soma}$: the effective radius of the soma
  • $S/V_{soma}$: the soma surface to volume ratio
Projections
  • $<R_{branch}>_{s}$: the mean effective radius of segments along branch s
  • $CV_{branch}$: a measure of branch beading
  • $L_{branch}$: the branch length
  • $S/V_{branch}$: the branch surface to volume ratio
  • $<\mu OD_{branch}>_{s}$: a measure of mean branch undulation
  • $<R_{c}>_{s}$: a measure mean branch curvature
  • $\tau_{branch}$: a measure of branch tortuosity
  • $\theta_{branch}$: the angle formed by two bifurcating branches
@@
  • $BO$: the degree of branching of the cellular projections
An illustration of the structural descriptors investigated is reported in Fig.2 and a summary of their definitions is reported in Tab.1.
@@
From the information in the cellular reconstruction SWC file, the nodes/edges defining the projections from those belonging to the soma were separated: all the nodes and corresponding edges within a distance 1.5x the nominal soma radius (radius of the first node in the SWC file) from the first node, namely the ’root’, were assigned to the soma; the remaining ones to the projections.
@@
Using this soma threshold the cell was resampled, preserving the nodes that lie within the soma threshold. From this resampled soma the 3D surface mesh was constructed using Blender and the volume, $V_{soma}$, and surface, $S_{soma}$, of the soma calculated. The soma volume can be expressed in terms of the effective soma radius $R_{soma}$, which is the radius of a sphere of equivalent volume. Additionally, we calculated also the soma surface-to-volume ratio, $S/V_{soma}$ as ratio between $S_{soma}$ and $V_{soma}$; and the fraction of the soma surface covered by the cellular projections as sum of the projection connection area divided by the soma surface area ($\eta_{soma}$).
@@
From the nodes/edges defining the projections, the individual branches composing the cellular projections were identified, delimited by either branching or termination nodes. The individual branches are comprised of cylindrical sub-segments, defined by the edges and their associated radius. The central line defined by these sub-segments gives a curvilinear path, $s$. From this path, $s$, metrics for the projections features were computed. Branch beading is reported as the coefficient of variation of the branch radius along the branch length ($CV_{branch}$). The branch surface-to-volume ratio ($S/V_{branch}$) was determined as the sum of the sub-segments area divided by the sum of the sub-segments volume. Branch undulation ($<\mu OD_{branch}>_{s}$) is calculated as the mean angle made by the vector of the individual sub-segments and the vector made by the branch start and end points.
@@
Finally, the general metrics were computed. The cellular domain $R_{domain}$ is defined as the maximum Euclidean distance from the root presented in each cell. The number of projections $N_{projection}$ was found by identifying the number of branches that cross the soma threshold. And the branch order $BO$ is defined as the number of consecutive bifurcations of the cellular projections.
@@
$R_{soma}$ Radius of sphere of volume equivalent to the soma surface volume
$<R_{branch}>_{s}(<R_{branch}^{2}>_{s})^{1/2}$ Mean (standard deviation) of subsegments’ radius along s
$CV_{branch}$ Coefficient of variation of radius
$L_{branch}$ Sum of subsegments’ length in s
$S/Z_{branch}$ Ratio between the sum of the surfaces of all the subsegments in s and the sum of their volume
$<\mu OD_{branch}>_{s}$ Mean microscopic orientation dispersion
$<R_{c}>_{s}$ Mean radius of curvature of s
@@
$BO$ The number of consecutive bifurcations of the cellular projections
$N_{proj}$ The number of primary projections radiating from the soma
$R_{domain}$ Distance of the furthest node from the soma
@@
We believe these characteristics can be used to model GM cells [47] and inform dMRI sequence design to maximise acquisition sensitivity to cellular features.
We provide the mean and standard deviation for each structural descriptor for each cell type for each species analysed. Moreover, we also provide value distributions for features of relevance to biophysical modelling of diffusion in GM, such as $R_{soma}$; $L_{branch}$ and estimated intracellular residence times based on the surface-to-volume ratio of the whole cell: $S/V_{domain}$ and the individual branches: $S/V_{branch}$.

### v4

@@ Abstract
Grey matter structure is a key focus in neuroscience, as cell morphology varies by type and can be affected by neurological conditions. Understanding these variations is essential for studying brain function and disease.
Diffusion-weighted MRI (dMRI) is a powerful tool for examining cellular microstructure in vivo, but its accuracy depends on identifying which morphological features influence its measurements. Despite growing interest, no systematic report has defined key neural cell traits.
We analysed more than 11,800 3D cellular reconstructions across three species and nine cell types, establishing reference values for critical traits. These fall into three categories: structural, shape, and topological features.
Beyond defining these traits, we assess their relevance for dMRI, identifying which neural features it can be sensitive to. This work provides essential benchmarks for gray matter research, aiding in the interpretation of neuroimaging data and improving brain tissue models.
To complement the statistical analysis, we also provide high resolution 3D surface meshes representative of each cell type and species. These meshes are fully compatible with Monte Carlo simulators, offering a valuable resource for the modelling community.
@@ 1 Introduction
Despite the widespread use of biophysical models in diffusion-weighted Magnetic Resonance Imaging (dMRI) to infer cellular-scale structure, gray matter (GM) remains poorly understood due to a lack of ground-truth morphological data. Unlike white matter (WM), where axonal morphology is well characterized, GM cellular features—critical for accurate dMRI modeling—are rarely quantitatively analysed. This knowledge gap leads to overly simplistic or unfounded model assumptions, potentially reducing the reliability of microstructural imaging in GM. Here, we address this challenge directly by systematically characterizing the statistical distribution of key morphological features across different species, providing a much-needed empirical foundation for improving GM dMRI models.
@@
Grey matter is composed of a range of cells, mainly differentiated into neuronal and glial cells, featuring a plethora of morphological characteristics. Neurons are fundamental functional units of the nervous system, specialized in the transmission and integration of electrical and chemical signals within the brain. Supporting the neurons are the glial cells (e.g. astrocytes, microglia and oligodendrocytes) that are crucial in maintaining health and functionality.
First studied in depth by Ramon y Cajal [1], neuronal morphology offers insights into the complex structure and function of the brain.
Neural cells exhibit a remarkable diversity in shape and size, each adapted to its specific function [2]. They can be classified based on morphological, molecular, and physiological characteristics [3]. Both neurons and glial cells share a common structural organization, consisting of a central soma and branching projections, yet they vary significantly across brain regions [4, 5]. For instance, Purkinje cells in the cerebellum display intricate dendritic arborization, whereas granule cells in the cerebral cortex exhibit a much simpler morphology. Furthermore, neurons of the same cell type can exhibit significant morphological differences across brain regions [6], and even display substantial variability between cortical layers within the same region [7].
@@
The brain contains approximately 86 billion of these neural111neural should not be confused with neuronal. Neural means relating to the nervous system (hence referring to any cell type in the brain: from neurons to glia), while neuronal means relating to neurons. cells [8]. Cortical GM is composed of 10–40% cell bodies (soma) of neural cells; 40–75% neurites: neuronal dendrites, short‐range intra‐cortical axons, the stems of long‐range axons extending into the WM and glial cell projections which intermingle with each other to form a dense and complex network; 15–30% highly tortuous extra-cellular space (ECS); and 1–5% vasculature [9, 10, 11]. In adults, the glia to neuron ratio is 1.32/1.40 for males/females respectively. The proportion of glial cells (by cell count) was estimated to be 77% oligodendrocytes, 17% astrocytes and 6% microglia [12, 13].
[…]
The morphology of neurons and glia is revealed through staining techniques such as confocal or electron microscopy [16, 17, 18], which can image cellular structures with very high resolution (down to a few nanometers) [19].
@@
Currently, there are no methods to directly observe cellular microstructure in vivo, as its scale (measured in micrometers) exceeds the resolution of clinical MRI, which typically operates at the millimeter scale [20]. While non-invasive imaging techniques like MRI can provide insights, they cannot directly capture cellular-level details. With many neurological conditions, such as dementia [21], as well as aging [22], altering the brain structure on this cellular scale, there is a strong incentive to develop means of revealing the morphology of neural cells, in-vivo.
@@
Although dMRI has still millimeter-scale resolution, it is sensitive to micrometer-scale structures by measuring the diffusion of endogenous molecules (e.g., water). This makes it a promising technique for overcoming MRI’s resolution limits and characterizing the brain’s microstructure (i.e. its cellular-scale organization) in vivo. However, dMRI’s sensitivity to microstructure is indirect, and biophysical modelling of the brain tissue and the subsequent interpretation of the dMRI signal is essential to quantify histologically meaningful features of the cellular structure, and gain specificity to their changes. To this end, the microstructure imaging paradigm has been introduced over a decade ago [23]: the approach fits a biophysical model voxel-wise to the set of signals obtained from images acquired with different sensitisations to tissue microstructure, yielding maps of model parameters that it is hoped are proxies of the corresponding underlying microstructural features.
@@
Successful examples of the microstructure imaging paradigm include Neurite Orientation Dispersion and Density Imaging (NODDI) [24] and the White Matter Tract Integrity (WMTI) [25] to characterise the diffusion of water within WM, revealing insight into the structure of axonal bundle tracks and other anatomical features, such as axon diameter [26, 27].
@@
Building on the success of dMRI in WM, there has been growing interest in applying it to GM to characterize cellular morphology in vivo[28, 29, 30, 31, 32, 33, 34]. dMRI has already been used to distinguish different cortical regions and reveal laminar structures within GM [35]. Significant effort has been made to develop also models to better describe the diffusion signal within GM and reveal anatomical information about the microstructure, such as the Soma And Neurite Density Imaging (SANDI) [32] to characterize soma and neurite density, the Neurite EXchange Imaging (NEXI) [33] and the Standard Model with EXchange (SMEX) [34] to characterize the water permeative exchange between neurites and extracellular space, and combinations of these, such as SANDI with exchange (SANDIx) [34].
@@
While substantial effort has been made to design, validate, and translate to clinics biophysical models for WM [23, 36, 37], the GM counterpart is lagging. This disparity stems from the greater complexity of the tissue, which renders the design of biophysical models for GM microstructure imaging more challenging [38].
Building accurate dMRI models for GM requires a fundamental understanding of its microstructural features. Key questions remain: what morphological properties influence the dMRI signal? How can they be measured reliably? How should these properties guide model development? With cellular morphology varying by cell type as well as brain region, there is a necessity for a thorough analysis of characteristics based on these criteria.
@@
Here, we aim to correct this imbalance with a comprehensive analysis of GM cellular morphology, looking at structural and topological morphology, and shape descriptors of 11,850 real three-dimensional (3D) reconstructions from mouse, rat, monkey, and human brain cortex. To complement the statistical analysis, we provide 50 high resolution 3D surface meshes for each cell type and species; fully compatible with Monte Carlo simulators, offering a valuable resource for the modelling community.
@@
The paper is organized as follows. We first provide quantitative information on the anatomy of brain GM tissue at the cellular scale. We characterize the morphology of neural cells using structural, topological and shape descriptors. We then review the range of dMRI measurements and biophysical models available to probe this anatomy and highlight limitations and caveats, ultimately providing guidelines on how to model GM microstructure from dMRI signals.
2 Materials and Methods
@@ 2.1 Microscopy Dataset
In order to measure characteristic features for specific cell types, the open access repository Neuromorpho.org [39] which has a comprehensive range of cellular reconstructions available was used. We downloaded and analysed three dimensional reconstructions of 11,850 brain cells, from mouse, rat, monkey, and human, in the form of SWC files  [^2] The ”SWC” encodes for the last names of its initial designers Ed Stockley, Howard Wheal, and Robert Cannon and is an ASCII text-based file that describes three-dimensional neuronal or glial morphology.. The SWC file defines a set of labelled nodes connected by edges characterizing the three-dimensional structure of each cell Fig.1.
Nine representative cell-types were acquired: microglia, astrocyte, oligodendrocyte, pyramidal, granule, purkinje, glutamatergic, gabaergic, and basket cells, from mouse/rat (N=9,001), monkey (N=525), and human (N=2,324). Although the NeuroMorpho database contains over 270,000 cellular reconstructions, only 11,850 satisfied our inclusion criteria: healthy controls; having complete reconstruction of soma and all the dendrites; containing information about diameters and angles; and being 3D reconstructions (see Supplementary Fig.1). Our quality assessment criteria include: consistent estimates of dendritic diameters (e.g., most of the rejected reconstructions had a fixed nominal diameter for all the branches instead of the real one); continuity of the cellular processes.
Despite their essential role in synaptic formation and the brains micro connectivity, dendritic spines (small protrusions on neuronal dendrites that form synaptic connections) were largely absent from the reconstructions. Furthermore, their inclusion biases the statistical analysis of overall cellular morphology, such as artificially reducing branch length and inflating branch order. As a result, for all reconstructions spines were identified (if present) and removed. Additionally, due to the inconsistency in axonal reconstructions (axons largely being absent or heavily truncated) in the reconstructions, they were not considered in the analysis (if axonal components were present, they were identified from the swc format and removed from the reconstruction). As a result, the statistics of the projections are for dendritic projections only.
The cellular reconstructions were analysed in Matlab using custom scripts, exploiting functions from validated suites (TREES [40], Blender [41], ToolboxGraph [42]). All the codes and SWC files used in this work will be made publicly available on GitHub https://github.com/Charlie-Aird/Decoding-Grey-Matter upon publication.
@@ 2.2 Structural descriptors
The structural analysis describes the constituent parts of the cell, including the soma and cellular projections, offering vital information about the cell’s fundamental structure, such as the effective soma radius and the branch angle between daughter branches.
For the structural analysis, such features or descriptors, determined to be crucial to microstructure modelling based on current literature [43, 32, 44, 45, 46, 47, 48, 49, 50, 51, 52], were estimated from the acquired cellular reconstructions. This set of features allows for a deeper understanding of how each fundamental aspect of brain cell morphology influences the diffusion of molecules within the intracellular space. For instance, the size of the soma and cellular projections can provide insights into the characteristic length scales of intracellular restrictions, while the branching, tortuosity, undulation, and calibre variation of the projections can inform on time-dependent diffusion processes. Additionally, the surface-to-volume ratios of the soma and projections offer valuable information on exchange dynamics.
The features of the soma and cellular projections differ significantly, and our analysis accounts for this by organizing the structural descriptors into relevant categories. These categories are: (1)soma (the characteristics defining the cell body); (2)projections (the set of characteristics defining the cellular projections’ structure as interconnected branches) and (3)general (describing the general cellular characteristics).
@@
  • BOBO: the degree of branching of the cellular projections
  • $S/V_{domain}$: the surface-to-volume ratio of the complete cellular structure (soma and branches)
Soma
  • $R_{soma}$: the effective radius of the soma (the radius of a sphere of equivalent volume as the soma)
  • $R_{MRsoma}$: the effective MR radius of the soma radii distribution
  • $S/V_{soma}$: the soma surface-to-volume ratio
Projections
  • $R_{branch}>_{s}$: the mean effective radius of segments along branch s
  • $R_{MRbranch}>_{s}$: the effective MR radius of the branch radii distribution
  • $CV_{branch}$: a measure of branch beading
  • $L_{branch}$: the branch length
  • $S/V_{branch}$: the branch surface-to-volume ratio
  • $\mu OD_{branch}>_{s}$: a measure of mean branch undulation
  • $R_{c}>_{s}$: a measure mean branch curvature
  • $\tau_{branch}$: a measure of branch tortuosity
  • $\theta_{branch}$: the angle formed by two bifurcating branches
Fig.2 illustrates these descriptors, and a summary of their definitions is reported in Tab.1.
@@
From the information in the cellular reconstruction SWC file, the nodes/edges defining the projections from those belonging to the soma were separated: all the nodes and corresponding edges within the nominal soma radius (radius of the first node in the SWC file) from the first node, namely the ’root’, were assigned to the soma; the remaining ones to the projections.
@@
Using this soma threshold the cell was resampled, preserving the nodes that lie within the soma threshold. From this resampled soma the 3D surface mesh was constructed using Blender and the volume, $V_{soma}$, and surface, $S_{soma}$, of the soma calculated. The soma volume can be expressed in terms of the effective soma radius $R_{soma}$.
From the distribution of soma radii the effective MR soma radius was calculated using the following equation [34],
$R_{MRsoma}=(<R_{soma}^{7}>/<R_{soma}^{3}>)^{(1/4)}$
Additionally, we calculated also the soma surface-to-volume ratio ratio as $S_{soma}/V_{soma}$; and the fraction of the soma surface covered by the cellular projections as sum of the projection connection area divided by the soma surface area ($\eta_{soma}$).
@@
From the nodes/edges defining the projections, the individual branches composing the cellular projections were identified, delimited by either branching or termination nodes. The individual branches are comprised of cylindrical sub-segments, defined by the edges and their associated radius. The central line defined by these sub-segments defines a curvilinear path, $s$. From this path, $s$, metrics for the projections features were computed. Branch beading is reported as the coefficient of variation of the branch radius along the branch length ($CV_{branch}$). The branch surface-to-volume ratio ($S/V_{branch}$) was determined as the sum of the sub-segments area divided by the sum of the sub-segments volume. Branch undulation ($\mu OD_{branch}>_{s}$) is calculated as the mean angle subtended by the vector of the individual sub-segments and the vector made by the branch start and end points.
Branch radius, $R_{branch}$, is reported as the average branch radius along path $s$. From the distribution of branch radii the effective MR branch radius was calculated using equation [27],
$R_{MRbranch}=(<R_{branch}^{6}>/<R_{branch}^{2}>)^{(1/4)}$
@@
Finally, the general metrics were computed. The number of projections $N_{projection}$ was found by identifying the number of branches that cross the soma threshold. And the branch order $BO$ is defined as the number of consecutive bifurcations of the cellular projections.
@@
$R_{domain}$ Distance of the furthest node from the soma
$N_{proj}$ The number of primary projections radiating from the soma
BOBO The number of consecutive bifurcations of the cellular projections
$S/V_{domain}$ Ratio between the complete (soma and branches) cellular surface and cellular volume
$R_{soma}$ Radius of sphere of volume equivalent to the soma volume
$R_{MRsoma}$ Effective MR soma radius [34]
$S/V_{soma}$ Ratio between the soma mesh surface and soma mesh volume
$S/V_{branch}$ Ratio between the sum of the surfaces of all the subsegments in s and the sum of their volume
$L_{branch}>$ Mean of the Sum of subsegments’ length in s
$R_{branch}>_{s}(<R_{branch}^{2}>_{s})^{1/2}$ Mean (standard deviation) of subsegments’ radius along s
$R_{MRbranch}$ Effective MR branch radius $(<R_{soma}^{6}>/<R_{soma}^{2})^{(1/4)}>$ [34]
$CV_{branch}$ Coefficient of variation of branch radius
$\mu OD_{branch}>_{s}$ Mean microscopic orientation dispersion of subsegments along s
$R_{c}>_{s}$ Mean radius of curvature of s
@@
We provide the first, second, and third quartiles for each structural descriptor for each cell type for each species analysed. Moreover, we provide value distributions for features of relevance to biophysical modeling of diffusion in GM, such as $R_{soma}$$L_{branch}$, $S/V_{branch}$ and $S/V_{domain}$.
We also estimated intracellular residence times, $\tau_{ic}$, based on the surface-to-volume ratio of the whole cell ($S/V_{domain}$) and the individual branches ($S/V_{branch}$) using the following equation [53]:
$\tau_{i}=\frac{1}{S/V\kappa}$
where $\kappa$ is the membrane permeability, and corresponding exchange times [54]:
$\tau_{ex}=\tau_{i}f_{ec}$
where $f_{ec}$ is the extracellular volume fraction, assumed to be $30\%$.
Additionally, the relationship between structural features was analysed by calculating the Spearman’s rank correlation coefficient.